CAVITY-ENHANCED DIRECT FREQUENCY COMB VELOCITY MODULATIO SPECTROSCOPY

\underline{LAURA SINCLAIR} , KEVIN COSSEL, WILLIAM AMES, JUN YE AND ERIC CORNELL, JILA, University of Colorado Boulder, and the National Institute for Standards and Technology, Boulder, Colorado 80309..

We have developed a novel technique for broad bandwidth and high resolution survey spectroscopy of molecular ions. Cavity-enhanced direct frequency comb spectroscopy (CE-DFCS) provides broad bandwidth and high resolution by using individual comb lines as parallel detection channels. Here we combine CE-DFCS with velocity modulation spectroscopy to provide ion-specific detection with further enhanced sensitivity. The first application of this technique will map the electronic states of HfF^{+} and ThF^{+}, which are integral to the JILA electron electric dipole moment experiment.